OpenStudy (anonymous):

Francisco has a savings account balance of $2,033.88. The interest rate on the account is 2.9% compounded monthly. If he opened the account nine years ago, what was the value of his initial deposit? (2 points) Select one: a. $701.34 b. $932.18 c. $1,567.15 d. $1,975.81

OpenStudy (anonymous):

The formula for compount interest is: \[A=P(1+i)^n\] We know his present amount, \(A\), and we know \(n\) and \(i\) so we can isolate for \(P\): \(A=$2033.88\) \(i=\frac{0.029}{12}\) \(n=9*12\) \(2033.88=P(1+\frac{0.029}{12})^{108}\) \(2033.88=1.297819P\) \(P\dot{=}1567.15\) Therefore the answer is C